OpenStudy (anonymous):

find six trig functions if tan(x)=√3/3 and cos(x)=-√3/2

OpenStudy (anonymous):

i can not understand ! can u explain what do you want

OpenStudy (anonymous):

if you mean how to know the angel for example its (7*pi)/6

OpenStudy (zpupster):

sin = y/r csc=r/y cos = x/r sec=r/x tan = y/x cot= x/y tan = sqrt 3/3 = y/x you know your y and x cos = -sqrt3/2 you know your y and x and r fill in the rest